A hypothetical closed economy has a national income model of the form y=C+I+G where C=30+0.8Y and I and G and private Investment expenditure are exogenously determined to= 50 and 80 units respectively. Compute the national equilibrium level of income for this economy using aggregate income equal aggregate expenditure and withdrawals and equal injection methods
1
Expert's answer
2020-07-30T11:09:57-0400

1.Using aggregate income equal aggregate expenditure method:

"Y=C+I+G"

"Y=30+0.8Y+50+80"

"Y-0.8Y=30+50+80"

"0.2Y=160"

"Y=\frac{160}{0.2}"


"Y=800"


2.Using withdrawals and equal injection method:

"Value \; of \; withdrawals=Value \; of\; injections"

"Savings=Investment+Government \; expenditure"

"Y-C=I+G"

"Y-30-0.8Y=50+80"

"Y-0.8Y=50+80+30"

"0.2Y=160"

"Y=\frac{160}{0.2}"


"Y=800"